It is well known that bullets and other missiles fired at Superman simply bounce off his chest. Suppose that a gangster sprays Superman's chest with 7 g bullets at the rate of 105 bullets/min, the speed of each bullet being 450 m/s. Suppose too that the bullets rebound straight back with no change in speed. What is the magnitude of the average force on Superman's chest from the stream of bullets?
N​
Answer #1

Average Force on the chest = rate of change of momentum (p)

initial velocity of bullet = 450 m/s

Initial momentum = 450 × 7*10^-3

Final velocity = - 450 m/s

Final momentum = -450 ×7*10^-3

Force exerted by each bullet = 7*10^-3 (450 - (-450))

= 6.3 N

Number of bullet hitting per second = 105/60 = 1.75

Average force = 6.3 × 1.75 = 11 N